### v2

@@ Abstract
Rock weathering is a common phenomenon to consider in various engineering applications that involve underground operations, such as underground storage or geothermal energy extraction. This weathering is induced by the circulation of a reacting fluid in the pores of the rock and can induce changes in materials properties, but also on the stress state. This work presents discrete element simulations of sedimentary rocks dissolution to study their stress state evolution. The rock is modeled as a cohesive granular material subjected to debonding. Oedometric conditions are considered during the weathering and the evolution of the coefficient of lateral earth pressure $k_{0}$, a proxy of the stress state, is tracked. In particular, the influence of the degree of cementation, the confining pressure, the initial value of $k_{0}$ and the history of loading are investigated. It has been observed that cemented granular materials tend to reach an attractor configuration for the stress state with the increase of the degree of dissolution. $k_{0}$ aims to reach the $k_{0}^{attr}$ value, between 0.3 and 0.4, when all the bounds are dissolved independently of the initial state. Two main mechanisms have been observed to explain this evolution of the interparticles force configurations and thus of the stress state: the collapse of the unstable chain forces (stable only due to the cementation) and the softening of the grains.
@@
[…]
[1]organization=Institute of Mechanics, Materials and Civil Engineering, UCLouvain, addressline=Place du Levant 1, city=Louvain-la-Neuve, postcode=1348, country=Belgium
[…]
[2]organization=Multiphysics Geomechanics Lab, Duke University, addressline=Hudson Hall Annex, Room No. 053A, city=Durham, postcode=27708, country=NC, USA
@@
The influence of several parameters on the evolution of the ratio of horizontal to vertical stress, i.e. the $k_{0}$ coefficient, is illustrated.
@@ 1 Introduction
Most of the rocks are affected by weathering, especially in engineering applications like underground storage [1] or geothermal energy [2], which involve the modification of pore water chemistry or the circulation of a reacting fluid in the pores. In particular, underground storage strategies envisioned to store large amounts of $CO_{2}$, captured from large point sources like power generation facilities, or $H_{2}$, produced during overproduction periods of renewable energies like solar or wind, involve the injection of a fluid into a reservoir rock. This fluid reacts with the surrounding rock inducing dissolution and/or precipitation that affects rocks’ permeability [3], but also their mechanical behavior [4] or their stress state [5]. To ensure the long-term success and safety of such storage solutions, the role of chemo-mechanical couplings needs to be understood and modeled as they can influence settlement at the surface [6, 7], the stability of the caprock [8, 9] or induce seismicity [10].
The influence of weathering on the mechanical behavior of geomaterials has already been investigated experimentally [11, 12, 13, 14] and numerically [15, 16, 17, 18] for rocks (bonded material) and sands (unbonded material) in different configurations, but only a few studies have looked into the influence of weathering on the stress state of the material.
@@
A seminal paper on the evolution of the stress state during weathering was published by Vaughan and Kwan in 1984 [19], in which they developed a theoretical model to study the evolution of the stress state during the weathering of rocks to form residual soils. More recently, experimental investigations of dissolution have been conducted by soft oedometer tests, which allows the measurement of radial stresses, using calcarenite and artificially cemented sand [20], glass beads mixed with salt grains [21] or carbonate sand [22]. Most of those studies have focused on non-cohesive granular materials, except for [20], in which they observe for both the artificially cemented silica sand and the fully calcareous soft rock an increase of the radial stress during dissolution by an acid solution, albeit at different rates, until an asymptotic value is reached.
@@
This theoretical and experimental work has been completed by some numerical studies using mostly the discrete element method. Again, to our knowledge all existing numerical studies focus on non-cohesive granular materials and the dissolution is modeled as a homogeneous decrease of the diameter. In [21] they were able to reproduce experimental results, where they first observe a decrease of $k_{0}$ reaching the Rankine active pressure coefficient $k_{a}$ during dissolution, and $k_{0}$ may eventually increase to recover its initial value if dissolution continues. In [22], they have investigated the effect of dissolution kinetics and have observed the same behavior as in [21] for fast dissolution (a decrease of $k_{0}$ followed by an increase) and a monotonic decrease when the dissolution is slow.
The work presented in the present paper aims at investigating the evolution of $k_{0}$ with numerical simulations solved by discrete element modelization for rock materials. The rock material is described as a cohesive granular material following the model introduced in [23]. The mechanical parameters calibrated by Sarkis et al. [24] on biocemented sands [25] have been considered.
@@
Several parameters such as the degree of cementation, the confining pressure, and the history of load are investigated here. Especially, the initial value of the coefficient of lateral earth pressure, a proxy of the state of stress, is monitored and studied. Indeed, this parameter can be affected by tectonic solicitations in the context of underground reservoirs [26, 27, 28] and reaches a larger value than in experimental setups.
@@
The first section of this paper is dedicated to the formulation of the model used in the discrete element modelization. Then, a second one describes the framework and the context of this work. Finally, results and discussions are proposed in the next sections.
@@ 2 Theory and formulation
The Discrete Element Model (DEM) is an approach developed by Cundall & Strack [29] to simulate granular materials at the particle level and was extended to rock materials by [30]. The foundation of this method is to consider inside the material the individual particles and their interactions explicitly [31]. Newton’s laws (linear and angular momentum) are used to compute the motion of the grains, formulated as follows for one grain:
@@
where $m$ is the particle mass, $v_{i}$ is the particle velocity vector, $g_{i}$ is the gravity acceleration vector, $f_{i}$ is the sum of contact force vectors applied to the particle, $I$ is the moment of inertia of the particle, $\omega_{i}$ is the angular velocity vector, $M_{i}$ is the sum of contact moment vectors applied to the particle (torques due to bending, and to twisting and to the tangential forces).
@@
where $v^{12}_{i}$ is the relative velocity vector between grains and $\delta=\Delta_{ni}n^{12}_{i}$ is the norm of the normal overlap vector. Notice that the terms $\left(R-\delta/2\right)$ represent the corrected radii at the contact. Here, the angular velocity vectors $\omega_{i}$ of the grains are considered for computing the tangential overlap vector $\Delta_{si}$. As the contact orientation can evolve with time, it is important to update by rotation and scaling the tangential overlap vector $\Delta_{si}^{new}=\Delta_{si}^{old}-\Delta_{sj}^{old}n_{j}^{12}n_{i}^{12}$ and $\overline{\Delta_{si}^{new}}=\overline{\Delta_{si}^{old}}$.
@@
Those relative angular velocity vectors $\Delta\omega_{ti}$ and $\Delta\omega_{bi}$ are used to compute a twisting and a bending relative angular rotation vectors $\Delta\theta_{ti}$ and $\Delta\theta_{bi}$ incrementally.
@@
As the contact orientation can evolve with time, it is important to update by rotation and scaling the twisting and bending relative angular rotation vectors $\Delta\theta_{ki}^{new}=\Delta\theta_{ki}^{old}-\Delta\theta_{kj}^{old}n_{j}^{12}n_{i}^{12}$ and $\overline{\Delta\theta_{ki}^{new}}=\overline{\Delta\theta_{ki}^{old}}$, with $k=t\text{ or }b$.
The contact models between particles considered here obey a cohesive law following [23]. Normal, tangential, bending and twisting models are shown in Figure 1 and described in the following.
@@
The normal force vector $F_{ni}$ is described in Equation 13. An elastic stiffness $K_{n}$ is needed, formulated Equation 14.
@@
where $E_{m}$ is the macroscale Young’s modulus, $R^{1}$ and $R^{2}$ are the radii of the particles in stake.
The tangential force vector $F_{si}$ is described in Equation 15. An elastic stiffness $K_{s}$ is needed, formulated Equation 16.
@@
where $\nu$ is the macroscale Poisson’s ratio.
The bending moment vector $M_{bi}$ is described in Equation 17. An elastic stiffness $K_{b}$ is needed, formulated Equation 18.
@@
where $\alpha_{b}$ is a non-dimensional factor, linking the bending and the tangential stiffnesses.
The twisting moment vector $M_{ti}$ is described in Equation 19. An elastic stiffness $K_{t}$ is needed, formulated Equation 20.
@@
where $\alpha_{t}$ is a non-dimensional factor, linking the twisting and the tangential stiffnesses. The bending and the twisting resistances aim to reproduce the shape of the grain [32, 33, 34] as spheres are used in this framework (numerically more efficient).
Those contact laws are linear in the elastic domain. Some contact failure is introduced based on the Mohr-Coulomb theory presented in Equation 21.
@@
where $\mu$ is the friction coefficient between two particles.
The cementation of the sample is modeled as a cohesion between the grains. The bond exists until one of the two criteria presented in Equation 22 is not verified.
@@
where $\sigma_{s}$ is the shear strength of the bond, $\sigma_{n}$ is the tensile strength of the bond and $A_{b}$ is the surface of the bond.
@@ 3 Numerical model
The initial condition algorithm is presented in Figure 2. A box of $8~{}\times~{}2,8~{}\times~{}2,$ 8 × 2 , 8 mm is generated. Then $3000$ particles are created. It is important to notice that the grains are initially incorporated with a radius smaller than the final one, as a radius expansion algorithm is applied to generate an initial condition [31]. This algorithm aims to verify a uniform grain size distribution ($R_{min}=75$ $\mu$m and $R_{max}=125$ $\mu$m). Once the particles reach their final dimension, the position of the top wall is controlled to apply a vertical pressure equal to $P_{cementation}$ (see Equation 23 for details). Once the vertical pressure is equal to $P_{cementation}$, the friction between the grains, the twisting resistance and the bending resistance are switched on. Then, the cementation between the particles is applied. At every contact, there is a random draw (with a probability equal to $p_{c}$) to determine if a bond is generated. Those bonds verify a shear strength $\sigma_{s}$ and a tensile strength $\sigma_{n}$ and they have a surface $A_{b}$ obtained by a lognormal distribution presented in A and defined by the parameters $m_{log}$ and $s_{log}$ [24]. Once the bonds are generated, the position of the top wall is controlled to apply a vertical pressure equal to $P_{confinement}$ (see Equation 23 for details). It is important to notice that the sample is under oedometric conditions (the control of the top plate aims to verify at any instant the vertical confining pressure) [7, 20]. Then, the position of the lateral wall is controlled to reach an initial coefficient of lateral earth pressure $k_{0}=\frac{\sigma_{II}}{\sigma_{I}}$ ($\sigma_{I}$ is the vertical pressure, $\sigma_{II}$ is the lateral pressure). Once the lateral pressure is obtained, the position of the lateral plate is fixed to verify oedometric conditions (fixed lateral walls), and the control of this element is switched off. An example of an initial configuration is illustrated in Figure 3.
@@
Figure 2: Initial condition algorithm: a) A box is created. b) All the particles are generated with a reduced radius. c) The different radii increase steeply until reaching the final values. d) A vertical pressure is applied by moving the top plate of the box until verifying the vertical pressure $P_{cementation}$. e) The friction, the twisting resistance and the bending resistance are switched on at the contacts. f) The cementation is applied by generating bonds at the contacts. g) A vertical pressure is applied by moving the top plate of the box until verifying the vertical pressure $P_{confinement}$. h) A lateral pressure is applied by moving the lateral plate of the box until verifying the coefficient of lateral earth pressure $k_{0}$.
@@
Figure 3: An example of an initial configuration obtained by the algorithm presented in Figure 2. The sample is then in oedometric conditions: the lateral and bottom walls are fixed and a constant vertical pressure is applied at the top.
The control of the wall (top or lateral) is done by a proportional controller $k_{p}$ described in Equation 23. A maximum speed of the plate is applied to verify quasi-static conditions [31].
@@
where $v_{plate}$ is the velocity of the plate (in the direction of the control), $k_{p}$ is the value of the proportional controller, $P_{plate}$ is the pressure applied on the plate (in the direction of the control), $P_{target}$ is the targeted pressure applied on the plate (in the direction of the control) and $v_{plate}^{max}$ is the maximum velocity of the plate.
About the walls, no friction, no bending resistance and no twisting resistance are considered with the grains.
@@
while Left some bonds do
     Dissolve the bonds $\triangleright$ the bonds surfaces are decreased
@@
     while not DEM equilibrium do $\triangleright$ granular reorganization occurs
@@
Once the initial configuration is set, the dissolution steps start. As presented in Algorithm 1, those steps are divided into several parts. First, the bonds are dissolved. The surface of each bond not broken is reduced (the value of the decrease must be small enough to stay representative). Then, the reorganization of the grains occurs until reaching an equilibrium. This state is defined by two characteristics :
  • 1.
    the unbalanced force (defined as the ratio of the mean summary force on bodies and mean force magnitude on interactions) is smaller than a criteria value (here $0.01$).
  • 2.

### v3

@@ Abstract
Rock weathering is a common phenomenon to consider in various engineering applications that involve underground operations, such as underground storage or geothermal energy extraction. This weathering is induced by the circulation of a reacting fluid in the pores of the rock and can induce changes in materials properties and stress state. This work presents discrete elements simulations of the dissolution of sedimentary rocks to study their stress state evolution. The rock is modeled as a cohesive granular material subjected to debonding. Oedometric conditions are considered during the weathering and the evolution of the coefficient of lateral earth pressure $k_{0}$, a proxy of the stress state, is tracked. In particular, the influence of the degree of cementation, the confining pressure, the initial value of $k_{0}$ and the history of loading are investigated. It has been observed that cemented granular materials tend to reach an attractor configuration for the stress state with the increase of the degree of dissolution. $k_{0}$ aims to reach the $k_{0}^{attr}$ value, between 0.3 and 0.4, when all the bounds are dissolved, independently of the initial state. Two main mechanisms have been observed to explain this evolution of the interparticles force configurations and thus of the stress state: the collapse of the unstable chain forces (stable only due to the cementation) and the softening of the grains.
@@
  • The influence of several parameters on the evolution of the ratio of horizontal to vertical stresses, i.e. the $k_{0}$ coefficient, is illustrated.
@@ 1 Introduction
To diminish the impact of human activity on climate evolution, one approach involves the utilization of renewable energy sources, such as wind and solar power, to generate electricity. However, these energy sources are characterized by variability, necessitating the development of strategies for energy storage and utilization. A promising approach involves the transformation of excess energy into $H_{2}$. The challenge, however, lies in identifying suitable methods to store the produced $H_{2}$, which must be subsequently extracted and consumed during periods of energy deficiency. It has been posited that underground storage and salt caverns hold considerable promise for this purpose, offering a greater storage capacity than conventional aboveground infrastructure such as pipelines or tanks [1, 2]. Likewise, geothermal energy appears as an efficient method to produce energy, and it is defined as the injection of a fluid underground to extract the thermal energy of the Earth [3, 4]. Furthermore, deep aquifers have emerged as a promising medium for the sequestration of greenhouse gases, including $CO_{2}$ [5]. This engineering application, added to the emergence of renewable energy production, participates in the reduction of the impact of human activities on climate change. If $H_{2}$ storage and geothermal energy production involve a fluid injection and extraction, $CO_{2}$ sequestration implies only injection, necessitating underground retention. To assess the viability of the $H_{2}$ storage and $CO_{2}$ sequestration techniques, various pilot projects have been initiated [6, 7, 8, 9].
The injection of the fluid into a geomaterial reservoir induces a modification of the fluid chemical composition, and thus its reactivity with the surrounding solid phases, activating or enhancing chemical processes [10, 11]. The consequences are numerous: settlement at the surface [12, 13, 14, 15], modifications of the mechanical [2, 16, 17] and hydraulic [18, 19, 20] properties, or threat to the stability of the caprock [21, 22]. It has even been correlated to earthquake nucleation [23, 24, 25, 26]. These phenomena and consequences observed in the context of the previously described engineering applications are similar to the numerous examples in nature where geomechanics and chemistry appear deeply intertwined, such as weathering [27, 28]. Even if the kinetics of the natural processes is much slower than the kinetics of the engineering-induced consequences, Terzaghi pointed out that long-term phenomena are the main causes of the instabilities rather than single events [29], requiring investigation.
The impact of weathering, due to natural processes or engineering applications, on the mechanical behavior of geomaterials has been previously investigated through both experiments [30, 31, 32, 33] and numerical simulations [17, 34, 35, 36, 37, 38, 39, 40] for sands (unbonded material) and rocks (bonded material) in a range of configurations. However, only a limited number of studies have focused on examining the influence of weathering on the stress state of the material, even if it can induce the material to reach the yield limit [41].
@@
A seminal paper on the evolution of the stress state during weathering was published by Vaughan and Kwan in 1984 [28]. In this work, the authors developed a theoretical model to study the evolution of the stress state during the weathering of rocks to form residual soils. More recently, experimental investigations of dissolution have been conducted using soft oedometer tests, which allow the measurement of radial stresses [42]. These tests have been performed with calcarenite and artificially cemented sand [43], glass beads mixed with salt grains [44] or carbonate sand [45]. The majority of these studies have focused on non-cohesive granular materials, except [43], in which the authors observed an increase in radial stress during dissolution by an acid solution in both the artificially cemented silica sand and the fully calcareous soft rock. However, the rate of increase was different in each case, reaching an asymptotic value in both cases.
@@
The theoretical and experimental work has been completed through the implementation of numerical studies, predominantly using the Discrete Element Method [46, 47]. It should be noted that, to the best of our knowledge, all existing numerical studies focus on non-cohesive granular materials and model dissolution as a homogeneous decrease in diameter. In [44], the authors were able to reproduce experimental results in which an initial decrease of $k_{0}$ was observed. This coefficient reaches the Rankine active pressure coefficient $k_{a}$ during dissolution. Additionally, the authors noted that $k_{0}$ may eventually increase to recover its initial value if dissolution continues. In their investigation of the effect of dissolution kinetics [45], the authors have observed a similar behavior for rapid dissolution, a decrease followed by an increase of $k_{0}=\sigma_{II}/\sigma_{I}$, with $\sigma_{I}$ and $\sigma_{II}$ are the vertical and lateral stresses. However, they observed a monotonic decrease for slower dissolution rates.
The objective of the present study is to investigate the evolution of $k_{0}$ through numerical simulations solved by Discrete Element Modelization for rock materials. In the context of rocks, weathering can be conceptualized as three distinct phenomena depicted in Figure 1. The first chemical alteration is the substitution of a mineral for one with a weaker stiffness or lower density [48]. The second phenomenon is the dissolution of grains and the bonds located between them [43]. This degradation is exemplified by calcarenite, wherein the grains and bonds are composed of calcium carbonate. Indeed, the dissolution kinetics for the bond and the grain appear comparable. The third case is the dissolution of the bonds only, while the grains remain intact [43]. An illustrative example is provided by silicic sand, which contains carbonate bonds. The dissolution kinetics of the bond is assumed to be considerably faster than that of the grain. This phenomenon is referred to as debonding and is the subject of the present work. Ciantia et al. have even identified two distinct kinds of bonding within the rock, temporary and persistent bonding, affecting the weathering process [31]. The temporary bonding origin is identified as the mineral suspension which deposits if the material is dried. In the case of fluid saturation of the sample, these bonds dissolve, with a fast kinetics reaction, as the material is suspended in the pore fluid. Conversely, the persistent bonding origin is the diagenesis phenomenon, and the dissolution kinetics is much smaller. This differentiation between the bonds involves a short and long-term aspect of the mechanical evolution of the geomaterial during weathering.
Figure 1: Scheme of the weathering in the context of rock. Three phenomena are distinct: a) the mineral replacement by a weaker phase, b) the dissolution of grains and bonds (calcarenite for example), and c) the dissolution of the bonds, while the grains remain intact (silica sand + carbonate bonds for example).
Despite the potential of constitutive relations to address chemo-mechanical couplings, these approaches remain constrained to the macroscale. The phenomena occurring at the microscale, such as granular reorganization and material dissolution, are merely postulated rather than modeled. Furthermore, these formulations often consider a homogeneous dissolution within the sample, which is not in agreement with experiments where wormhole formations are observed while injecting acid into a triaxial cell [49, 50]. Similarly, chemo-mechanical responses appear localized, and not homogeneous, in the host rock during $CO_{2}$ sequestration [9]. The question that remains is how numerical models which explicitly consider the microstructure predict the mechanical behavior of geomaterials subjected to debonding.
The Discrete Element Model (DEM) [47], developed in 1979 by Cundall and Strack, has emerged as a pertinent method to predict the mechanical behavior of sands [46] and rocks [51] at the microscale. In the seminal approach, the granular material is modeled as spheres. The existing bonds in rock are conceptualized as additional stiffnesses between the grains. The rock material is described as a cohesive granular material following the model introduced in [52]. The mechanical parameters were calibrated by Sarkis et al. [53] on biocemented sands [54]. Indeed, these biocemented sands can be considered as a suitable proxy for natural sandstones [55, 56]. While the pressures considered here may not fully reflect the conditions present in an underground reservoir, the observed behavior can be readily extrapolated through an equivalent contact stiffness number ($\propto$ Young modulus/pressure applied) [57, 58]. This discrete description can be readily extended to capture the chemistry of the weathering. Indeed, the size of the bonds can be reduced to model the silica + carbonate case. These modifications destabilize the granular microstructure and affect the properties at the sample scale.
@@
The present study investigates several parameters, including the degree of cementation, the confining pressure, and the history of load. In particular, the initial value of the coefficient of lateral earth pressure, which serves as a proxy for the state of stress, is monitored and studied. Indeed, this parameter can be affected by tectonic solicitations in the context of underground reservoirs [59, 60, 61], reaching a larger value than in experimental setups.
@@
The first Section of this paper is dedicated to the formulation of the model utilized in the Discrete Element Modelization. Subsequently, a second Section describes the framework and the context of this work. Finally, results and discussions are proposed in the next Sections.
@@ 2 Theory and formulation
The Discrete Element Model (DEM) is an approach developed by Cundall & Strack [46] for simulating granular materials at the particle level. This model was extended to rock materials by [51], adding a bond between the particles. The fundamental premise of this methodology is to explicitly consider the individual particles and their interactions within the material itself [47]. Newton’s laws (linear and angular momentum) are employed to compute the motion of the grains, which is formulated as follows for a single grain:
@@
where $m$ is the particle mass, $v_{i}$ is the particle velocity vector, $g_{i}$ is the gravity acceleration vector, $f_{i}$ is the sum of contact force vectors applied to the particle, $I$ is the moment of inertia of the particle, $\omega_{i}$ is the angular velocity vector, $M_{i}$ is the sum of contact moment vectors applied to the particle (torques due to bending, twisting, and tangential forces). To solve these equations, various integration methods can be used [62]. For example, the software YADE used during this work follows a Verlet scheme [63]. The time integration depends on a time step $dt$ that must verify the P-wave critical time step condition [64] defined as:
$dt<dt_{crit}=\frac{\text{min}(R)}{\sqrt{E_{m}/\rho}}$ (3)
where $R$ is the radius of the grains, $E_{m}$ is the Young modulus of the contact and $\rho$ is the density of the material. Herein, a safety factor is considered, $dt=0.6\,dt_{crit}$.
@@
where $v^{12}_{i}$ is the relative velocity vector between grains and $\delta=\overline{\Delta_{ni}}$ is the norm of the normal overlap vector. It should be noted that the terms $\left(R-\delta/2\right)$ represent the corrected radii at the contact. Here, the angular velocity vectors $\omega_{i}$ of the grains are taken into account for the calculation of the tangential overlap vector $\Delta_{si}$. As the contact orientation can evolve, it is crucial to update the tangential overlap vector by rotation and scaling: $\Delta_{si}^{new}=\Delta_{si}^{old}-\Delta_{sj}^{old}n_{j}^{12}n_{i}^{12}$ and $\overline{\Delta_{si}^{new}}=\overline{\Delta_{si}^{old}}$.
@@
These relative angular velocity vectors $\Delta\omega_{ti}$ and $\Delta\omega_{bi}$ are used to compute a twisting and a bending relative angular rotation vectors $\Delta\theta_{ti}$ and $\Delta\theta_{bi}$ incrementally.
@@
As the contact orientation can evolve, it is crucial to update the twisting and bending relative angular rotation vectors by rotation and scaling: $\Delta\theta_{ki}^{new}=\Delta\theta_{ki}^{old}-\Delta\theta_{kj}^{old}n_{j}^{12}n_{i}^{12}$ and $\overline{\Delta\theta_{ki}^{new}}=\overline{\Delta\theta_{ki}^{old}}$, with $k=t\text{ or }b$.
The contact models between particles considered here are governed by a cohesive law following [52]. This cohesion permits the relative displacement between bonded particles. However, due to the presence of cementation, an additional stiffness is introduced. The normal, tangential, bending and twisting models are illustrated in Figure 2 and described in the following.
@@
As illustrated in Figure 3, a cemented contact can be classified into three distinct types: I) Frictional, II) Mixed/Cohesive (Frictional+Cemented), or III) Cemented [53]. In the case of type I, no cementation is present. In types II and III, the contact can be modeled as two springs in parallel (type II) or in series (type III). One represents the bond $K_{\cdot,b}$ and the other represents the unbonded grain-grain contact $K_{\cdot,u}$. In the following Sections, the assumption is made that the sample is composed of contacts type II, as depicted in Figure 2 where springs are represented in parallel.
The contact force vector $f_{i}$ applied to the particle, see Equation 1, is defined as the sum of the normal force vectors $F_{ni}$ and the tangential force vectors $F_{si}$ over the contacts including the particle. The normal force vector $F_{ni}$ is described in Equation 14. An elastic stiffness $K_{n}=K_{n,b}+K_{n,u}$ is required, formulated Equation 15.
@@
where $E_{m}$ is the contact Young modulus, $R^{1}$ and $R^{2}$ are the radii of the particles in stake.
Similarly, the tangential force vector $F_{si}$ is described in Equation 16. An elastic stiffness $K_{s}=K_{s,b}+K_{s,u}$ is required, formulated Equation 17.
@@
where $\nu$ is the contact Poisson’s ratio.
In the same vein, the contact moment vector $M_{i}$ applied to the particle, see Equation 2, is defined as the sum of the bending moment vectors $M_{bi}$, the torque due to the tangential force vectors $F_{si}$, and the twisting moment vector $M_{ti}$ over the contacts including the particle. The bending moment vector $M_{bi}$ is described in Equation 18. An elastic stiffness $K_{b}=K_{b,b}+K_{b,u}$ is required, formulated Equation 19.
@@
where $\alpha_{b}$ is a non-dimensional factor, relating the bending and the tangential stiffnesses.
Similarly, the twisting moment vector $M_{ti}$ is described in Equation 20. An elastic stiffness $K_{t}=K_{t,b}+K_{t,u}$ is required, formulated Equation 21.
@@
where $\alpha_{t}$ is a non-dimensional factor, relating the twisting and the tangential stiffnesses. The bending and the twisting resistances aim to reproduce the shape of the grain [65, 66, 67] as spheres are used in this framework (numerically more efficient).
These contact laws are linear in the elastic domain. Furthermore, a Coulomb friction limit is introduced at the contact level, see Equation 22
@@
where $\mu$ is the friction coefficient between two particles. This friction coefficient is constant before and after the bond breakage.
The bond exists until one of the two criteria presented in Equation 23 is not verified.
@@
where $\sigma_{s}$ is the shear strength of the bond, $\sigma_{n}$ is the tensile strength of the bond and $A_{b}$ is the surface of the bond. It is worth noting that no criterion is employed for the compression condition. This value of the bond surface $A_{b}$ is determined initially and is reduced during the simulation, see Section 3. During the debonding phenomenon, the mechanical rupture of the bond can be induced by a stress reorganization, increasing the tensile/shear force transmitted in the contact. Furthermore, the rupture can occur even if the tensile/shear force transmitted stays constant as the bond surface $A_{b}$ decreases during the dissolution. The last rupture case is due to chemistry when the bond surface reaches $A_{b}=0\,m^{2}$. Once the bond breaks (by mechanics or chemistry), the bond surface $A_{b}$ is set to $0\,m^{2}$ and the tensile stiffness of the contact is deactivated.
@@ 3 Numerical model
The Discrete Element Model is solved using the YADE open source software [63]. Some examples of scripts used are available on GitHub: https://github.com/AlexSacMorane/YADE_oedo_acid. The initial condition algorithm is presented in Figure 4. A $8~{}\times~{}2,8~{}\times~{}2,$ 8 × 2 , 8 mm box is generated. Subsequently, $3000$ particles are generated, despite $1600$ would suffice to obtain a representative elementary volume [47, 53]. It is crucial to note that the grains are initially incorporated with a radius smaller than the final one. Indeed, a radius expansion algorithm is applied to generate the initial condition [47]. The objective of this algorithm is to verify a uniform grain size distribution similar to [53] ($R_{min}=75$ $\mu$m and $R_{max}=125$ $\mu$m). Once the particles have reached their final dimension, the position of the top wall is controlled in order to apply a vertical pressure equal to $P_{cementation}$ (see Equation 24 for details). The positions of the remaining walls are maintained at a fixed position. Once the vertical pressure is equal to $P_{cementation}$, the friction between the grains, the twisting resistance, and the bending resistance are activated. The top wall is then controlled to attain once more a vertical pressure equal to $P_{cementation}$, while the other walls remain fixed. Subsequently, the cementation between the particles is applied. At each contact, a random draw is conducted with a probability equal to $p_{c}$ to ascertain whether a bond is formed. The bonds are defined by a shear strength $\sigma_{s}$ and a tensile strength $\sigma_{n}$. Additionally, the bonds have a surface area $A_{b}$ obtained by a lognormal distribution presented in Appendix A and defined by the parameters $m_{log}$ and $s_{log}$ [53]. The parameters $p_{c}$, $m_{log}$, and $s_{log}$ characterize the initial degree of cementation, see Table 2. Subsequently, the position of the top wall is controlled in order to apply a vertical pressure equal to $P_{confinement}$ (see Equation 24 for details). The remaining walls are maintained in a fixed position. Subsequently, the position of one lateral wall (the remaining ones are fixed) is controlled to reach an initial coefficient of lateral earth pressure $k_{0}={\sigma_{II}}/{\sigma_{I}}$ ($\sigma_{I}$ is the vertical pressure, $\sigma_{II}$ is the lateral pressure). The sample is subjected to oedometric conditions (the control of the top plate is active and it aims at verifying the vertical confining pressure at any given instant) [14, 43]. Once the lateral pressure has been obtained, the position of the moving lateral plate is fixed to verify the oedometric conditions (fixed lateral walls), and the control of this element is deactivated. The initial configuration is reached, and the destabilization due to the debonding phenomenon starts. An illustrative example of an initial configuration is provided in Figure 5.
@@
Figure 4: Initial condition algorithm: a) A box is created. b) All the particles are generated with a reduced radius. c) The different radii increase steeply until the final values are reached. d) A vertical pressure $P_{cementation}$ is applied by moving the top plate. e) The friction, the twisting resistance, and the bending resistance are activated. f) The cementation is applied by generating bonds at the contacts. g) A vertical pressure $P_{confinement}$ is applied by moving the top plate. h) An initial stress state $k_{0}$ is applied by moving the lateral plate.
a) Refer to caption   b) Refer to caption
@@
Figure 5: a) An example of an initial configuration obtained by the algorithm presented in Figure 4. b) The sample is then in oedometric conditions: the lateral and bottom walls are fixed and a constant vertical pressure is applied at the top during the injection of the fluid that induces debonding.
Bond breakage can be triggered during the initialization of the test, preceding the commencement of the bond dissolution through acid injection. The portion of the broken bond is given Table 5 in B. Bond breakage will inevitably occur as a consequence of the initialization algorithm, as illustrated in Figure 4. The cementation occurs at a specified pressure $P_{cementation}$, and the pressure is then elevated to the confining pressure $P_{confinement}$ before the start of the test. This increase is the source of some bond breakage (even prior to the commencement of dissolution). Moreover, the objective is to reproduce numerically the experiments. In an oedometric test with acid injection (bond dissolution), the sample is initially cemented, placed in the oedometer, loaded at confining pressure, and then subjected to acid injection (bond dissolution) [43]. In the experiments, the step of loading to the confining pressure also induces bond breakage. The occurrence of bond breakage is more prevalent in samples with $k_{0}^{init}\leq k_{0}^{attr}$. Indeed, the value of $k_{0}$ is approximately equal to $k_{0}^{attr}$ prior to the formation of the bonds. In order to achieve the objective of $k_{0}^{init}\leq k_{0}^{attr}$, it is necessary to extend the lateral dimension of the sample by moving the lateral wall. The combination of this extension with the application of the confining pressure at the top results in shear and tensile conditions at the contact levels. Consequently, bonds breakage occurs if one of the conditions presented in Equation 23 is reached. This phenomenon occurs less frequently in the case of $k_{0}^{init}\geq k_{0}^{attr}$. Indeed, the lateral dimension of the sample is compressed, and it results in less shearing and tensile solicitations at the contact level, but in compression of the bonds. This phenomenon induces less bond breakage since no compression condition is considered as a breakage criterion.
The control of the wall (in the vertical or lateral direction) is achieved through the use of a proportional controller $k_{p}$ described in Equation 24. A maximum speed of the plate $v_{plate}^{max}\left(=0.00005\,R_{mean}/\,dt\right)$ is applied to verify quasi-static conditions [47], where $R_{mean}$ is the mean radius and $dt$ is the time step defined Equation 3. These aforementioned parameters were determined through a trial-and-error method, minimizing the oscillation of the obtained confining pressure around the target value and maximizing the control kinetics.
@@
where $v_{plate}$ is the velocity of the plate (in the direction of the control), $k_{p}$ is the value of the proportional controller, $P_{plate}$ is the pressure applied on the plate (in the direction of the control), $P_{target}$ is the targeted pressure applied on the plate (in the direction of the control) and $v_{plate}^{max}$ is the maximum velocity of the plate. The walls are considered to have no friction, cohesion, bending resistance, and twisting resistance with the grains.
@@
while bonds are remaining do
     Dissolve the bonds $\triangleright$ the bonds’ surfaces are decreased
@@
     while DEM equilibrium not reached do $\triangleright$ granular reorganization occurs
@@
Once the initial configuration has been established, the dissolution process commences. As presented in Algorithm 1, these steps are subdivided into distinct phases. Initially, the bonds are dissolved, with the surface area of each intact bond reduced (the magnitude of the reduction must be sufficiently minor to ensure representativeness). Subsequently, the reorganization of the grains occurs until an equilibrium state is reached. This mechanical equilibrium is determined by two indices:
  • the unbalanced force (defined as the ratio of the mean summary force on bodies and mean force magnitude on interactions) is smaller than a criterion value (here $0.01$).